OpenStudy (anonymous):

how do I solve : CSC pi/6

OpenStudy (amistre64):

with a calculator?

OpenStudy (anonymous):

without please :D!

OpenStudy (amistre64):

memory then

OpenStudy (amistre64):

csc is the invert of sin

OpenStudy (amistre64):

and pi/6 is 30 degrees

OpenStudy (amistre64):

so invert the sine of 30 degrees

OpenStudy (anonymous):

then is it 2?

OpenStudy (amistre64):

lol ... it might be :) yep

OpenStudy (anonymous):

Sine of 30 deg is 1/2, so csc30 = 2

OpenStudy (anonymous):

great 2 more questions coming up. lol check the question plz

OpenStudy (anonymous):

find \[\frac{\pi}{6}\] on the last page of the cheat sheet. first coordinate is cosine, second coordinate is sine, and cosecant is the reciprocal of sine